OpenStudy (anonymous):

3 times the difference of x and y is -30, and 2/3 of x plus 1/2 of y is 19. what is x and y

OpenStudy (anonymous):

say what now?

OpenStudy (anonymous):

The question as i understand it is:\[3(x-y)=-30\]\[2/3x+1/2y=19\] I think I would solve by substituting (sorry about the working out) *** Trying to check if my answer is correct - one moment

OpenStudy (anonymous):

sorry, wrong thing (COPY AND PASTE)

OpenStudy (anonymous):

OK got it!\[3(x−y)=−30\]\[2/3x+1/2y=19\] I wanted to solve for x or y in ONE of the two equations. My choice was y. I started with the first equation: \[3(x−y)=−30\]Making x equal to 0 in the first equation we get \[-3y=-30\] That soon becomes\[y=10\] as we place -3 on the right hande side -30/-3 =30/3 We have the value of y. In the second equation\[2/3x+1/2y=19 \]But \[y=10\]So\[2/3x+1/2y=19 \]becomes \[2/3x+1/2(10)=19 \]which gives \[2/3x+5=19 \] Simplifying to give \[2/3x=14 \] So \[x=14\div(2/3)\] The answer being\[y=10\] and \[x=21\]

OpenStudy (anonymous):

O_O gee thanks. Wow